OpenStudy (anonymous):

solve for x 7(x + 3) + 3( x + 3)= 5x - 5

OpenStudy (anonymous):

multiply everything out, and get 7x+21 +3x+9 and combine like terms to get 10x+30 = 5x-5. Divide by 5 both sides. 2x+6 = x-1. put the x's on the same side. x+6=-1. Move the 6 over. X=-7

OpenStudy (anonymous):

7(x + 3) + 3( x + 3)= 5x - 5 7x+21+3x+9=5x-5 7x+3x-5x=-5-21-9 5x=-35 x=-7

OpenStudy (anonymous):